OpenStudy (anonymous):

13r^-4(r^-5)^5/8(r^2)^-2 solve with steps please.

OpenStudy (anonymous):

\[\frac{ 13r ^{-4}(r ^{-5})^{5} }{ 8(r ^{2})^{-2} }\]

OpenStudy (anonymous):

\[\frac{ 13*\frac{ 1 }{ r ^{4} } (\frac{ 1 }{ r ^{5} })^{5} }{ 8*\frac{ 1 }{ (r ^{2})^{2} } }\] Do you know how to do that?

OpenStudy (anonymous):

Not really... can you explain please?

OpenStudy (anonymous):

\[r ^{-100} = \frac{ 1 }{ r ^{100} }\] Is that clear?

OpenStudy (anonymous):

\[r ^{-4} = \frac{ 1 }{ r ^{4} }\]

OpenStudy (anonymous):

So now you have to simplify it \[\frac{ 13 *\frac{ 1 }{ r ^{4} * r ^{25} } }{ 8 * \frac{ 1 }{ r ^{4} } }\] 25 because 5*5 = 25 \[\frac{ 13*\frac{ 1 }{ r ^{29} } }{ 8 * \frac{ 1 }{ r ^{4} } }\] \[\frac{ \frac{ 13 }{ r ^{29} } }{ \frac{ 8 }{ r ^{4} } }\] (very up number is 13, hard to read) \[\frac{ 13 * r ^{4} }{ 8 * r ^{29} }\] 29-4=25 \[\frac{ 13 }{ 8r ^{25} }\] That should be resoult :)

OpenStudy (anonymous):

ah, perfect! thank you so much i was struggling with problems like these thanks got it :)

OpenStudy (anonymous):

Nice to hear that! Good luck!